What is the frequency of a photon whose energy is 6.00 x 10^-15
1
Expert's answer
2012-11-16T07:12:19-0500
Well, if the equation is E=hv, which also means energy of the quantum of radiation (E)= planck's constant (h) X frequency of the radiation emitted(v)"

so this is what it would look like:

6.00*10^-15 = 6.626*10^-34 X v

.906*10^19 = v

9.06*10^18 = the frequency (v)